OpenStudy (ray1998):

I can't, for the life of me, figure this out. Please help! Fan and medal will be given. Thank you in advance! :) Solve. 2(x-5) 3/2 = 54

OpenStudy (ray1998):

Yes :)

OpenStudy (mathmale):

Is that (3/2) an exponent or a multiplicand? I have my doubts because the expression as posted above includes (2/2) = 1. Double check this, please.

OpenStudy (ray1998):

It looks like an exponent. And I have written it exactly as the problem is given.

OpenStudy (ray1998):

Yes :)

OpenStudy (ray1998):

Is that how it starts out?

OpenStudy (anonymous):

2(x-5) ^(3/2) = 54 (x-5) ^(3/2) = 27 ((x-5) ^(3/2) )^(2/3) = 27^(2/3) x-5 = 9 x = 14